OpenStudy (anonymous):

A model M3SWE mandolin has a list price of $1299.90 with a chain discount of 5/3/2 a) what is the trade discount amount b)what is the net price?

OpenStudy (anonymous):

I got The Trade discount is 1299.9 x .95 = 1234.90

OpenStudy (anonymous):

and net price I got $400 x 1/5 = $80 400 - 80 $320 x 1/10 = $32 320 - 32 $288 x 1/20 = $14.40 288 - 14.40 $273.60 net price - Answer

OpenStudy (anonymous):

so what is 11.97.85 ? mean

OpenStudy (anonymous):

you can't take 1234.90/11.97.85